OpenStudy (anonymous):

Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold.

OpenStudy (anonymous):

you have two points (400,0) ,(0,300) first slope = y2−y1x2−x1=0−300400−0=−34

OpenStudy (anonymous):

then a line with slope m and passing point (x1,y1) is y-y1=m(x-x1) y−y1=m(x−x1)(400,0)m=−34y−0=−34(x−400)

OpenStudy (anonymous):

Okay. First find y

OpenStudy (anonymous):

next you need to find the slope of the line 2x+3y = 1470 so 3y = -2x + 1470 then divide by 3 and you get y=−23+490

OpenStudy (anonymous):

Hope this helped... have a great day!

OpenStudy (anonymous):

how did you get 3y = -2x + 1470